OpenStudy (anonymous):

Find the sum of the first 8 terms of the sequence. Show all work for full credit. 1, -3, -7, -11, ...

OpenStudy (anonymous):

this is an arithmatic sequence with \(a_0=1\) and a common difference of \(d=-3-1=-4\) sum of first n terms: \[S_n={n\over2}[2a_0+(n-1)d]\]

OpenStudy (anonymous):

haha for full credit, lol we aint getting any full credit.

OpenStudy (anonymous):

@some_someone lol I know

OpenStudy (anonymous):

@Ambition is that why its "ambition", Ambition?

OpenStudy (anonymous):

@electrokid lol yes, But for your equation, do I plug in the numbers?

OpenStudy (anonymous):

\[S _{n} = \frac{ n }{ 2 } (a _{1} + a _{n})\] where: \[a _{1} \] is the first term. \[a _{n} \] is the last term.